OpenStudy (cccmpt2):

A swimming pool is 16 meters wide and 25 meters long. At $50 per square meter, how much will a child-proof covering for the pool cost?

OpenStudy (cccmpt2):

I get $4,100. But, my study guide answer sheet says it's $20,000 ?????

OpenStudy (zpupster):

Area = length x width so 25 x 16 = 400 sq m At 50$ per sq m $50 x 400 = 20,000 $/sq m

OpenStudy (cccmpt2):

oh, i see! i used the formula for perimeter! this question concerns the space covering the pool, not surrounding it. thank you so much for your help! ~knocking self in forehead now!~

OpenStudy (whpalmer4):

Here's a tip that would have saved you: always use the units in your calculations! If you did the perimeter = 2*l + 2*w you would get 16 m * 2 + 25 m * 2 = 82 m * $50/m^2 = 4100 $/m $/m? Have you ever seen a piece of money with $/m on it? Hmm...something isn't right here... On the other hand, with the correct calculation: A = l*w = 16m*25m = 400 m^2 * $50/m^2 = $20,000 I haven't seen a $20,000 bill, but I've seen plenty of smaller bills with $ on them.